OpenStudy (anonymous):

Solving literal equations.. Can someone explain in simple terms how to do it? I know that it's just an equation with multiple variables, but I'm confused with the fractions they're giving me.. For example, they used s=1/2gt^2 and solved for g. They did some multiplication, then the fractions screwed me up, so can someone help me understand it? Thanks

OpenStudy (anonymous):

you have an equation with many variables. if they ask you solve for one of those, just do something to isolate that variable in 1 side, all of the others in other side. done. in your original equation, it is kind of " solve for s" so s =.....; if they ask you "solve for t, then t =....., " solve for g" then g =......... . you know what i mean?

OpenStudy (anonymous):

I don't give you the answer for your problem, just the way you have to do when meeting the question like that.

OpenStudy (anonymous):

I got that equation, but for S = 4v^2 solve for v, I divide both sides by 4, but then how do I isolate v? Divide by 1^2? That won't work, will it?

OpenStudy (goformit100):

I can Explain it to you.

OpenStudy (anonymous):

Well I think I get how to do them for the most part now, but I'm stuck on that other one now.. If there's just v squared, then how do I isolate it to get it without the square?

OpenStudy (anonymous):

\[ S = 4v^2\] \[\frac{S}{4}=v^2\] \[\pm\sqrt{\frac{S}{4}}=v\] \[v=\pm\frac{\sqrt{S}}{2}\]

OpenStudy (anonymous):

Thanks! That's what I was trying to figure out, the third step. I see what you did.

OpenStudy (anonymous):

yw